Three of six vertices of a regular hexagon are chosen at random. The probability that the triangle with these three vertices is equilateral equals

Grade:12

2 Answers

bharat bajaj IIT Delhi
10 years ago
Total number oftrianglesformed =6C3= 20
No. of equilateraltriangles= 2
? Required probability = 2 / 20 = 1 / 10
THanks
Bharat Bajaj
IIT Delhi
askiitians faculty
Shrey Joshi
7 years ago
student-name Japneet Kaur asked in Math3 of the 6 vertices of a regular hexagon are chosen at random find the probability that the triangle with three vertices is equilateral. 2 Follow 0student-name Ajanta answered this6807 helpful votes in Math, Class XII-Scienceout of 6 vertices, the number of ways to choose 3 vertices are the number of equilateral triangles that can be drawn from the vertices of a hexagon = 2 therefore the required probability = 2/2= = 1/10 hope this helps u